# SFMNet: Sparse Focal Modulation for 3D Object Detection

### v1

@@ 3.2 SFMNet
Building on our sparse focal modulation (SFM) module, we present SFMNet, a simple and efficient sparse detector for 3D object detection. It enables learning long-range dependencies that convolution-based backbones lack, while maintaining sparse convolution efficiency.
Our architecture is built on two core principles. First, increasing the effective receptive field (ERF) in mid- and high-level layers enhances the network’s capacity to capture broader spatial contexts. This is achieved by incorporating our proposed SFM module. Second, once a reasonable receptive field is achieved, scaling a model with small kernels becomes more efficient due to better parameter utilization. This aligns with [12]’s observation that beyond a certain point, adding large kernels is less efficient than adding small kernels. In our setup, SFM blocks can be seen as analogous to large kernels, while SRB blocks correspond to small kernels. Therefore, we include more SRB blocks than SFM blocks (see Section 4.2).
@@
Fig. 4(a) illustrates the SFMNet architecture. The input point cloud is voxelized, and features are extracted using a dynamic voxel feature extraction (VFE) method [64]. These features are then processed by our 3D and 2D sparse backbones. Our 3D backbone includes four stages and between each pair of consecutive stages, a regular sparse convolution with stride $2$ is used for downsampling. Each stage includes an SFM block followed by multiple submanifold residual blocks (SRBs) for high-level feature extraction (Fig. 4(b)). We use a variable number of SFM blocks and SRBs to enhance representational capacity and increase depth depending on the dataset. The higher the stage, the more SRB blocks per SFM block should be added to scale the model.
The SFM block utilizes our SFM module as the token mixer within a standard MetaFormer-like design [57], as shown in Fig. 4(c). Formally, given an input feature map $x\in\mathbb{R}^{N\times C}$ and its corresponding focal-modulated feature $z\in\mathbb{R}^{N\times C}$ generated by the SFM module, as described in Section 3.1, the SFM block is formulated as:
$y^{\prime}$ $=\text{LN}(z)+x$ (5)
$y$ $=\text{LN}(\text{MLP}(y^{\prime}))+y^{\prime},$ (6)
where LN denotes LayerNorm, and the multilayer perceptron (MLP) contains a single hidden layer. The 3D sparse features are then compressed into 2D sparse bird’s-eye view (BEV) features, following the approach in [58], and subsequently fed into a 2D backbone. Similar to the 3D backbone, the 2D backbone also consists of an SFM block followed by SRBs but employs 2D sparse convolutions instead of 3D.
@@ 4 Experiments
Datasets and metrics. To validate our approach’s effectiveness, we conducted experiments on three popular datasets: Argoverse2 [50], Waymo Open [41], and nuScenes [2]. The detection range of these datasets varies, with long-range detection in Argoverse2 (up to ±$\pm 200$ meters), mid-range in Waymo Open (up to ±$\pm 75$ meters), and short-range in nuScenes (up to ±$\pm 54$ meters). We use the standard evaluation metrics for each dataset: mAP for Argoverse2, mAP and mAP weighted by heading accuracy (mAPH) for Waymo Open, and mAP and nuScenes detection score (NDS) for nuScenes. On Waymo, the metrics are applied to two difficulty levels: L1 for objects with more than five LiDAR points and L2 for those with at least one sampled point.
@@
CenterPoint [56] 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
VoxelNeXt [7] 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
HEDNet [59] 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
SAFDNet [58] 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
SFMNet (Ours) 40.4 79.0 49.5 72.3 49.3 46.1 65.1 73.8 55.1 53.2 62.8 54.2 30.2 44.8 26.7 57.9 31.7 20.9 8.3 33.4 44.4 21.6 0.0 22.1 3.1 33.1 12.1
Table 2: Results on the Argoverse2 validation set (long-range, up to ±$\pm 200$ meters). The long-range detection dataset highlights our detector’s ability to capture long-range dependencies.
@@
Transformers SST [14] [CVPR’22] 74.5/71.0 67.8/64.6 74.2/73.8 65.5/65.1 78.7/69.6 70.0/61.7 70.7/69.6 68.0/66.9
TransFusion-L [1] [CVPR’22] -/- -/64.9 -/- -/65.1 -/- -/63.7 -/- -/65.9
SWFormer [43] [ECCV’22] -/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
CenterFormer [65] [ECCV’22] 75.6/73.2 71.4/69.1 75.0/74.4 69.9/69.4 78.0/72.4 73.1/67.7 73.8/72.7 71.3/70.2
DSVT-Voxel [47] [CVPR’23] 80.3/78.2 74.0/72.1 79.7/79.3 71.4/71.0 83.7/78.9 76.1/71.5 77.5/76.5 74.6/73.7
PTv3 [51] [CVPR’24] -/- 73.0/70.5 -/- 71.2/70.8 -/- 76.3/70.4 -/- 71.5/70.4
Large kernels LargeKernel3D [6] [CVPR’23] -/- -/- 78.1/77.6 69.8/69.4 -/- -/- -/- -/-
LinK [27] [CVPR’23] (6 epochs) -/- -/64.6 -/- -/65.0 -/- -/60.4 -/- -/68.4
Small kernels SECOND [52] [Sensors’18] 67.2/63.1 61.0/57.2 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0
PointPillars [23] [CVPR’19] 69.0/63.5 62.8/57.8 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9
Part-A2 [37] [TPAMI’20] 73.6/70.3 66.9/63.8 77.1/76.5 68.5/68.0 75.2/66.9 66.2/58.6 68.6/67.4 66.1/64.9
PV-RCNN [36] [CVPR’20] 76.2/73.6 69.6/67.2 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8
CenterPoint [56] [CVPR’21] 75.9/73.5 69.8/67.6 76.6/76.0 68.9/68.4 79.0/73.4 71.0/65.8 72.1/71.0 69.5/68.5
AFDetV2 [22] [AAAI’22] 77.2/74.8 71.0/68.8 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1
FSD [15] [NeurIPS’22] 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
PV-RCNN++ [38] [IJCV’23] 78.1/75.9 71.7/69.5 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2
VoxelNeXt [7] [CVPR’23] 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
HEDNet [59] [NeurIPS’23] 81.4/79.4 75.3/73.4 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9
SAFDNet [58] [CVPR’24] 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
@@
Table 3: Results on the Waymo Open validation set (mid-range, up to ±$\pm 75$ meters). Our SFMNet outperforms transformer- and large-kernel-based detectors, while achieving results comparable to the state-of-the-art small-kernel detector [58].
Implementations details. To compare with prior state-of-the-art methods, we trained SFMNet for $24$ epochs on Argoverse2 and Waymo Open, and $20$ epochs on nuScenes. Ablation studies used 6 epochs on Argoverse2. Further implementation details are in the supplemental materials.
@@ 4.1 Results
Quantitative results. We expect our model to be particularly beneficial for long-range scenes. Indeed, as shown in Table 2, which compares our method to state-of-the-art sparse convolution-based detectors with small kernels, our detector achieves a $0.7\%$ mAP improvement over the previous best method [58], demonstrating its ability to capture long-range dependencies. For mid-range detection on Waymo Open, Table 3 compares our method to state-of-the-art methods on the validation set, where ’-’ indicates unavailable results. Our detector outperforms both transformer-based and large kernel-based methods, achieving gains of $1.7\%$ and $3.3\%$ mAPH on Level-2 difficulty compared to the transformer-based detectors DSVT [47] and PTv3 [51], respectively, while also achieving comparable results to the best small kernel-based detector [58]. For the short-range detection dataset nuScenes,Table 4 shows that our method achieves SoTA results on the mAP metric and ranks second on the NDS metric, outperforming both large kernel and transformer-based methods. The effective receptive field (ERF) is $3.3$ meters for Argoverse2 and $2$ meters for Waymo Open and nuScenes.
@@
Figure 5: Effective receptive field (ERF) of SFMNet. (a) A typical input scene with height color-coded in shades of gray. (b,c) The ERF for each ground-truth object, color-coded from blue (outside ERF) to red (high attention regions). Red and orange rectangles highlight zoomed-in views of two regions. In the right (orange) zoom-in, four stationary cars and a van, plus two moving cars and a van, are shown. Without the SFM module (b), only a local portion of each vehicle is within the ERF. With the SFM module (c), the entire object is captured, providing more context and spatial understanding for accurate 3D detection.
Qualitative results. To qualitatively validate our detector’s ability to capture an extensive ERF, for each ground-truth object, we calculate the ERF by sampling a query voxel and analyzing its reach. Fig. 5 visualizes the ERF (color-coded from blue to red) of our detector with and without our SFM module. The more yellowish the color, the larger the ERF. For each ground-truth object, we randomly select one voxel to serve as the query and calculate the ERF associated with it. As intended, the dependency coverage is regional, resulting in large regional receptive fields. Without the SFM module (b), the detector’s receptive field is limited to local areas around each query, capturing only fragments of the objects. For instance, in the left zoom-in (red rectangle) in Fig. 5(b), there are four vehicles. The receptive field covers only portions of each vehicle, leaving most voxels outside this field (blue). In contrast, with our SFM module (c), the detector gains a broader receptive field, allowing it to capture even partially occluded vehicles. This is reflected in the four vehicles being highlighted in green, yellow, and red colors.
Transformers TransFusion-L [1] [CVPR’22] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
DSVT-Voxel [47] [CVPR’23] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.8 70.9
Large kernels LargeKernel3D [6] [CVPR’23] 69.1 63.9 85.1 60.1 72.6 41.4 24.3 85.6 70.8 59.2 72.3 67.7
LinK [27] [CVPR’23] 69.5 63.6 - - - - - - - - - -
Small kernels CenterPoint [56] [CVPR’21] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
VoxelNeXt [7] [CVPR’23] 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
PillarNeXt-L [24] [CVPR’23] 68.4 62.2 85.0 57.4 67.6 35.6 20.6 86.8 68.6 53.1 77.3 69.7
HEDNet [59] [NeurIPS’23] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
SAFDNet [58] [CVPR’24] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
Table 4: Results on the nuScenes validation set (short-range, up to ±$\pm 54$ meters). Our SFMNet outperforms all transformer- and large-kernel-based detectors, surpassing the previous SoTA small-kernel detector on the mAP metric and ranking second on the NDS metric.
4.2 Ablation study
Table 5 compares our SFM module to traditional global and local attention, as well as set attention [47], on the Argoverse2 dataset over six epochs. For this evaluation, we replaced an attention block in the 3D and 2D backbones with either a single SFM module or an alternative attention mechanism, demonstrating that our SFM achieves higher accuracy and efficiency. While global attention attains a high ERF, it quickly reaches memory limits ($80$ GB) due to the large voxel count. In contrast, local and set attention operate within a limited context window, with set attention outperforming local attention. However, our SFM module surpasses both in accuracy (mAP) and efficiency (memory usage) despite using a smaller context window.
Method Window size(m) Memory(GB) mAP
@@ 4.2 Ablation study
Table 5: Comparison to attention mechanism. We replace our SFM module with various attention mechanisms and report mAP and per-GPU memory peak usage. While global attention covers the entire ERF, it quickly exceeds memory limits. In contrast, Local-attention and Set-attention [47] restrict their focus to voxels within a local context window. Our SFM module outperforms both in accuracy and efficiency, even with a smaller context window.
Table 6 examines the effect of varying the number of SFM blocks and SRBs. It shows that increasing the number of SRBs after the SFM block in Stage 4 (A) improves accuracy with a modest parameter increase. Similarly, adding more SFM block repetitions with SRBs (B) boosts accuracy but at a higher parameter cost. A balanced approach, with moderate repetitions and additional SRBs (C), achieves a good trade-off between accuracy and parameters.
More ablation studies on SFM module placement and ERF size choices are provided in the supplemental materials.
@@
Table 6: Effect of SFM blocks and SRBs. Increasing SRBs after each SFM block in Stage 4 and the 2D backbone boosts accuracy. Adding more SFM block repetitions (Multiplier) further improves accuracy, though at the cost of more parameters. A balanced trade-off between accuracy and parameters is achieved in (C).
Limitations. SFMNet relies on sparse convolutions with custom CUDA operations, making it less deployment-friendly than other deep learning frameworks (e.g., PyTorch).

### v2

@@ 3.2 SFMNet
Building on our sparse focal modulation (SFM) module, we present SFMNet, a simple and efficient sparse 3D object detector. It enables learning long-range dependencies that convolution-based backbones lack, while maintaining sparse convolution efficiency.
Our architecture is built on two core principles. First, increasing the effective receptive field (ERF) in mid- and high-level layers enhances the network’s ability to capture broader spatial context—this is primarily achieved using a single SFM block. Second, once a sufficient receptive field is established, adding depth with SRB blocks becomes more efficient, as they act like small kernels and offer better parameter utilization. In practice, we find that one SFM block combined with one SRB block is sufficient to meet our design goals. When further deepening the network, we prioritize adding more SRB blocks rather than additional SFM blocks (see Section 4.2). This design choice aligns with the observation from [ding2024unireplknet] that, after the receptive field is adequately covered, small kernels become more efficient than large ones—an analogy that maps well to our use of SFM (large kernel equivalent) and SRB (small kernel equivalent) blocks.
@@
Fig. 4(a) illustrates the SFMNet architecture. The input point cloud is voxelized, and features are extracted using a dynamic voxel feature extraction (VFE) method [zhou2020end]. These features are then processed by our 3D and 2D sparse backbones. Our 3D backbone includes four stages and between each pair of consecutive stages, a regular sparse convolution with stride $2$ is used for downsampling. Each stage includes an SFM block followed by multiple submanifold residual blocks (SRBs) for high-level feature extraction (Fig. 4(b)). We use a variable number of SFM blocks and SRBs to enhance representational capacity and increase depth depending on the dataset. The higher the stage, the more SRB blocks per SFM block should be added to scale the model.
The SFM block employs our SFM module as the token mixer within a standard MetaFormer-like design [yu2022metaformer], as illustrated in Fig. 4(c), where non-empty voxels serve as the tokens. Formally, given an input feature map $x\in\mathbb{R}^{N\times C}$ and its corresponding focal-modulated feature $z\in\mathbb{R}^{N\times C}$ generated by the SFM module, as described in Section 3.1, the SFM block is formulated as:
$y^{\prime}=\text{LN}(z)+x,\quad y=\text{LN}(\text{MLP}(y^{\prime}))+y^{\prime},$ (5)
where LN denotes LayerNorm, the multilayer perceptron (MLP) has a single hidden layer, and $y$ is the refined feature from focal sparse modulation. The 3D sparse features are then compressed into 2D bird’s-eye view (BEV) features, as in [zhang2024safdnet], and passed to a 2D backbone. Like the 3D backbone, the 2D backbone includes an SFM block followed by SRBs, but uses 2D sparse convolutions.
@@ 4 Experiments
Datasets and metrics. To validate our approach’s effectiveness, we conducted experiments on three popular datasets: Argoverse2 [wilson2023argoverse], Waymo Open [waymo], and nuScenes [nuscenes]. The detection range of these datasets varies, with long-range detection in Argoverse2 (up to $\pm 200$ meters), mid-range in Waymo Open (up to $\pm 75$ meters), and short-range in nuScenes (up to $\pm 54$ meters). We use the standard evaluation metrics for each dataset: mAP for Argoverse2, mAP and mAP weighted by heading accuracy (mAPH) for Waymo Open, and mAP and nuScenes detection score (NDS) for nuScenes. On Waymo, the metrics are applied to two difficulty levels: L1 for objects with more than five LiDAR points and L2 for those with at least one sampled point.
@@
CenterPoint 22.0 67.6 38.9 46.5 16.9 37.4 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 6.3 3.9 0.5 20.1 22.1 0.0 3.9 0.5 10.9 4.2
VoxelNeXt 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
HEDNet 37.1 78.2 47.7 67.6 46.4 45.9 56.9 67.0 48.7 46.5 58.2 47.5 23.3 40.9 27.5 46.8 27.9 20.6 6.9 27.2 38.7 21.6 0.0 30.7 9.5 28.5 8.7
SAFDNet 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
SFMNet 40.4 79.0 49.5 72.3 49.3 46.1 65.1 73.8 55.1 53.2 62.8 54.2 30.2 44.8 26.7 57.9 31.7 20.9 8.3 33.4 44.4 21.6 0.0 22.1 3.1 33.1 12.1
Table 1: Results on the Argoverse2 validation set (long-range, up to $\pm 200$ meters). The long-range detection dataset highlights our detector’s ability to capture long-range dependencies.
@@
Transformers SST [fan2022embracing] 74.5/71.0 67.8/64.6 74.2/73.8 65.5/65.1 78.7/69.6 70.0/61.7 70.7/69.6 68.0/66.9
TransFusion-L [bai2022transfusion] -/- -/64.9 -/- -/65.1 -/- -/63.7 -/- -/65.9
SWFormer [sun2022swformer] -/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
CenterFormer [zhou2022centerformer] 75.6/73.2 71.4/69.1 75.0/74.4 69.9/69.4 78.0/72.4 73.1/67.7 73.8/72.7 71.3/70.2
DSVT-Voxel [wang2023dsvt] 80.3/78.2 74.0/72.1 79.7/79.3 71.4/71.0 83.7/78.9 76.1/71.5 77.5/76.5 74.6/73.7
PTv3 [wu2024point] -/- 73.0/70.5 -/- 71.2/70.8 -/- 76.3/70.4 -/- 71.5/70.4
Large kernels LargeKernel3D [chen2023largekernel3d] -/- -/- 78.1/77.6 69.8/69.4 -/- -/- -/- -/-
LinK [lu2023link] (6 epochs) -/- -/64.6 -/- -/65.0 -/- -/60.4 -/- -/68.4
Small kernels SECOND [second] 67.2/63.1 61.0/57.2 72.3/71.7 63.9/63.3 68.7/58.2 60.7/51.3 60.6/59.3 58.3/57.0
PointPillars [pointpillars] 69.0/63.5 62.8/57.8 72.1/71.5 63.6/63.1 70.6/56.7 62.8/50.3 64.4/62.3 61.9/59.9
Part-A2 [parta2] 73.6/70.3 66.9/63.8 77.1/76.5 68.5/68.0 75.2/66.9 66.2/58.6 68.6/67.4 66.1/64.9
PV-RCNN [pvrcnn] 76.2/73.6 69.6/67.2 78.0/77.5 69.4/69.0 79.2/73.0 70.4/64.7 71.5/70.3 69.0/67.8
CenterPoint [yin2021center] 75.9/73.5 69.8/67.6 76.6/76.0 68.9/68.4 79.0/73.4 71.0/65.8 72.1/71.0 69.5/68.5
AFDetV2 [hu2022afdetv2] 77.2/74.8 71.0/68.8 77.6/77.1 69.7/69.2 80.2/74.6 72.2/67.0 73.7/72.7 71.0/70.1
FSD [fan2022fully] 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
PV-RCNN++ [pvrcnn++] 78.1/75.9 71.7/69.5 79.3/78.8 70.6/70.2 81.3/76.3 73.2/68.0 73.7/72.7 71.2/70.2
VoxelNeXt [chen2023voxelnext] 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
HEDNet [zhang2024hednet] 81.4/79.4 75.3/73.4 81.1/80.6 73.2/72.7 84.4/80.0 76.8/72.6 78.7/77.7 75.8/74.9
SAFDNet [zhang2024safdnet] 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
@@
Table 2: Results on the Waymo Open validation set (mid-range, up to $\pm 75$ meters). Our SFMNet outperforms transformer- and large-kernel-based detectors, while achieving results comparable to the state-of-the-art small-kernel detector [zhang2024safdnet].
Transformers TransFusion-L [bai2022transfusion] 70.1 65.5 86.9 60.8 73.1 43.4 25.2 87.5 72.9 57.3 77.2 70.3
DSVT-Voxel [wang2023dsvt] 71.1 66.4 87.4 62.6 75.9 42.1 25.3 88.2 74.8 58.7 77.8 70.9
Large kernels LargeKernel3D [chen2023largekernel3d] 69.1 63.9 85.1 60.1 72.6 41.4 24.3 85.6 70.8 59.2 72.3 67.7
LinK [lu2023link] 69.5 63.6 - - - - - - - - - -
Small kernels CenterPoint [yin2021center] 66.5 59.2 84.9 57.4 70.7 38.1 16.9 85.1 59.0 42.0 69.8 68.3
VoxelNeXt [chen2023voxelnext] 66.7 60.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
PillarNeXt-L [li2023pillarnext] 68.4 62.2 85.0 57.4 67.6 35.6 20.6 86.8 68.6 53.1 77.3 69.7
HEDNet [zhang2024hednet] 71.4 66.7 87.7 60.6 77.8 50.7 28.9 87.1 74.3 56.8 76.3 66.9
SAFDNet [zhang2024safdnet] 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
Table 3: Results on the nuScenes validation set (short-range, up to $\pm 54$ meters). Our SFMNet outperforms all transformer- and large-kernel-based detectors, surpassing the previous SoTA small-kernel detector on the mAP metric and ranking second on the NDS metric.
Implementations details. To compare with prior state-of-the-art methods, we trained SFMNet for $24$ epochs on Argoverse2 and Waymo Open, and $20$ epochs on nuScenes. Ablation studies used 6 epochs on Argoverse2. Additional implementation details are provided in the supplemental.
@@ 4.1 Results
Quantitative results. We expect our model to be particularly beneficial for long-range scenes. Indeed, as shown in Table 3, which compares our method to state-of-the-art sparse convolution-based detectors with small kernels, our detector achieves a $0.7\%$ mAP improvement over the previous best method [zhang2024safdnet], demonstrating its ability to capture long-range dependencies. For mid-range detection on Waymo Open, Table 3 compares our method to state-of-the-art methods on the validation set, where ’-’ indicates unavailable results. Our detector outperforms both transformer-based and large kernel-based methods, achieving gains of $1.7\%$ and $3.3\%$ mAPH on Level-2 difficulty compared to the transformer-based detectors DSVT [wang2023dsvt] and PTv3 [wu2024point], respectively, while also achieving comparable results to the best small kernel-based detector [zhang2024safdnet]. For the short-range detection dataset nuScenes,Table 3 shows that our method achieves SoTA results on the mAP metric and ranks second on the NDS metric, outperforming both large kernel and transformer-based methods. The effective receptive field (ERF) is $3.3$ meters for Argoverse2 and $2$ meters for Waymo Open and nuScenes.
@@
Figure 5: Effective receptive field (ERF) of SFMNet. (a) A typical input scene with height is shown in grayscale. (b,c) The ERF for each ground-truth object is color-coded from blue (outside ERF) to red (high-attention regions). Red and orange rectangles highlight zoomed-in regions. In the right (orange) zoom-in, four stationary cars and a van, plus two moving cars and a van, are shown. Without the SFM module (b), only a local portion of each vehicle is within the ERF. With SFM (c), the entire object is captured, providing more context and spatial understanding for accurate 3D detection.
Qualitative results. To qualitatively validate our detector’s ability to capture an extensive ERF, for each ground-truth object, we calculate the ERF by sampling a query voxel and analyzing its reach. Fig. 5 visualizes the ERF (color-coded from blue to red) of our detector with and without our SFM module. The more yellowish the color, the larger the ERF. For each ground-truth object, we randomly select one voxel to serve as the query and calculate the ERF associated with it. Indeed, the dependency coverage leads to large regional receptive fields. Without the SFM module (b), the receptive field is limited to local areas around each query, capturing only fragments of the objects. For instance, in the left zoom-in in Fig. 5(b), there are four vehicles. The receptive field covers only portions of each vehicle, leaving most voxels outside this field (blue). In contrast, with our SFM module (c), the detector gains a broader receptive field, allowing it to capture even partially occluded vehicles. This is reflected in the four vehicles being highlighted in green, yellow, and red colors.
Method Window(m) Memory(GB) mAP
@@
Table 4: Comparison to attention. While global attention covers the entire ERF, it quickly exceeds memory limits. In contrast, Local-attention and Set-attention [wang2023dsvt] focus on voxels within a local context window. Our SFM module outperforms both in accuracy and efficiency, even with a smaller window.
@@
Table 5: Effect of SFM blocks & SRBs. Stacking SRBs after each SFM block improves accuracy (A). Increasing SFM repetitions boosts it further (B), but with more parameters. Configuration (C) offers a balanced trade-off.
4.2 Ablation study
Table 4 compares our SFM module to traditional global and local attention, as well as set attention [wang2023dsvt], on the Argoverse2 dataset over six epochs. For this evaluation, we replaced an attention block in the 3D and 2D backbones with either a single SFM module or an alternative attention mechanism, demonstrating that our SFM achieves higher accuracy and efficiency. While global attention attains a high ERF, it quickly reaches memory limits ($80$ GB) due to the large voxel count. In contrast, local and set attention operate within a limited context window, with set attention outperforming local attention. However, our SFM module surpasses both in accuracy (mAP) and efficiency (memory usage) despite using a smaller context window.
Table 5 examines the effect of varying the number of SFM blocks and SRBs. Increasing the number of SRBs after the SFM block in Stage 4 (A) improves accuracy with a modest parameter increase. Similarly, adding more SFM block repetitions (Multipliers) with SRBs (B) boosts accuracy but at a higher parameter cost. A balanced approach, with moderate repetitions and additional SRBs (C), provides a good trade-off between accuracy and parameters.
#levels dilation rates ERF(m) Memory(GB) mAP
2 [1,3] 0.9 30.19 35.7
3 [1,3,5] 1.9 32.76 36.8
3 [1,5,9] 3.1 32.98 36.4
4 [1,3,5,7] 3.3 37.53 37.1
5 [1,3,5,7,9] 5.1 36.39 37.1
Table 6: ERF effects. We evaluate different ERF sizes by increasing the number of context levels, each using a sparse convolution with kernel size $3$ and varying dilation. Four levels, yielding an ERF of $3.3$ meters, achieve the best accuracy.
Table 6 investigates the impact of different context levels with varying ERF. Increasing levels improves accuracy but also raises memory usage. Beyond a point, more levels do not help, as the network tends to ignore overly long-range features. Relying only on high dilation rates without local contexts (third row) yields suboptimal performance, highlighting the need for short-range dependencies. Additional ablations on SFM placement are in the supplemental.
Limitations. Our method is designed to capture long-range dependencies and enhance models that lack this capability. On datasets requiring long-range reasoning, it improves not only those models but also others with a similar objective. However, on short-range data, while it still benefits the former, its performance is on par with the latter.